A youth group with 26 members is going skiing. Each of the five chaperones will drive a van or sedan. The vans can seat seven people, and the sedans can seat five people. Assuming there are no empty seats, how many of each type of vehicle could transport all 31 people to the ski area in one trip?
Answer by checkley77(12844)

x+y=5 or x=5-y
7x+5y=31
7(5-y)+5y=31
35-7y+5y=31
-2y=31-35
-2y=-4
y=-4/-2
y=2 sedans are used.
x=5-2=3 vans are used.
Proof:
7*3+5*2=31
21+10=31
31=31